OpenStudy (anonymous):

Please helpp!!!!!!!! Based on the graph of the following system of equations, determine the number of solutions. 2x + 3y = 9 6y = 5 − 4x No solutions Exactly one solution Exactly two solutions Infinitely many solutions

OpenStudy (anonymous):

do you know matrices?

OpenStudy (anonymous):

no :(

OpenStudy (whpalmer4):

@chris00 what part of "based on the graph" isn't clear to you? :-)

OpenStudy (anonymous):

"another way of checking"

OpenStudy (anonymous):

lets first see wat algebra tells us then we can graph

OpenStudy (whpalmer4):

doesn't have anything to check yet :-)

OpenStudy (anonymous):

;P

OpenStudy (anonymous):

graphing is too easy

OpenStudy (anonymous):

wait I want the easy way :o lol

OpenStudy (anonymous):

hahaha You want to learn new things! haah

OpenStudy (anonymous):

the fastest way at least I have 10 problems left on a 80 question study guideIm so tired

OpenStudy (whpalmer4):

Here's the easy way to graph it. Substitute 0 for x in first equation, find y. Substitute 0 for y in first equation, find x. Now you've got two points (the x and y intercepts). Draw a line through them. Repeat with the other equation.

OpenStudy (anonymous):

alright from the equations, we can easy see that one is a scalar multiple of 2 but there constants are different, hence the lines are parallel and distinct. thus no solutions.

OpenStudy (anonymous):

don't really need to graph to be honest.

OpenStudy (anonymous):

thats the easiest way

OpenStudy (anonymous):

Wait so whpalmer4 I have to draw a whole graph out?

OpenStudy (whpalmer4):

Don't be scared by "draw a graph" — you just have to plot 4 points and draw 2 straight lines. It's not like you're graphing some function where you have to plot a zillion points to see all the wiggles

OpenStudy (anonymous):

Okay can you help me plug 0 in like type it while I graph please?

OpenStudy (whpalmer4):

They're trying to get you a visual for the concept. Two lines that intersect represent a system with one solution: the point of intersection. Two lines that are parallel never cross obviously, so have no solutions. Two lines that are coincident have a solution everywhere, so infinitely many.

OpenStudy (anonymous):

Yes

OpenStudy (whpalmer4):

2x+3y=9 plug in 0 for x: 2(0) + 3y= 9 y = 3 so plot point (0,3) plug in 0 for y: 2x + 3(0) = 9 2x=9 x=4.5 so plot point (4.5,0) draw a line through them. that's your first line.

OpenStudy (anonymous):

your nomenclature is off a bit. you can have lines that are parallel and coincident and lines that are parallel and distinct.

OpenStudy (whpalmer4):

http://mathworld.wolfram.com/CoincidentLines.html

OpenStudy (anonymous):

okay so i have my first line and lol we are in court im judge judy

OpenStudy (anonymous):

lol u believe a website over a maths professor. lmao

OpenStudy (anonymous):

and chis is the defendant

OpenStudy (whpalmer4):

coincident is a more restrictive description than parallel, but yes, I'm implying "parallel" to mean not also coincident Have you met the chap who wrote that article?

OpenStudy (anonymous):

@Loser66 thanks for the message

OpenStudy (anonymous):

lol your being defensive mate. chill.

OpenStudy (anonymous):

And thank you whpalmer4 and chris for the help :)

OpenStudy (anonymous):

no worries but.

OpenStudy (anonymous):

bud*

OpenStudy (anonymous):

well peace guys and XD could I have a medal for response

OpenStudy (whpalmer4):

other points are (0,5/6) and (5/4, 0)

OpenStudy (anonymous):

and so its no solution right?

OpenStudy (anonymous):

u bet

OpenStudy (anonymous):

cool thanks yo XD

OpenStudy (anonymous):

thanks for the medal too im going to fan you

OpenStudy (anonymous):

good luck with your study!